### v1

# sample.bib

Abstract
With the development of artificial intelligence and breakthroughs in deep learning, large-scale Foundation Models (FMs), such as GPT, CLIP, etc., have achieved remarkable results in many fields including natural language processing and computer vision. The application of FMs in autonomous driving holds considerable promise. For example, they can contribute to enhance scene understanding and reasoning. By pre-training on rich linguistic and visual data, FMs can understand and interpret various elements in a driving scene, and provide cognitive reasoning to give linguistic and action commands for driving decisions and planning. Furthermore, FMs can augment data based on its understanding of driving scenarios to provide feasible scenes of those rare occurrences in the long tail distribution that are unlikely to be encountered during routine driving and data collection. The enhancement can subsequently lead to the improvement in the accuracy and reliability of autonomous driving systems. Another testament to the potential of FMs applications lies in the development of World Models, exemplified by the DREAMER series, which showcase the ability to comprehend physical laws and dynamics. Learning from massive data under the paradigm of self-supervised learning, World Model can generate unseen yet plausible driving environment, facilitating the enhancement in the prediction of road users behavior and the off-line training of driving strategies. In this paper, we synthesize the applications and future trends of FMs in autonomous driving. By utilizing the powerful capabilities of FMs, we strive to tackle the potential issues stemming from the long-tail distribution in autonomous driving, consequently advancing overall safety in this domain.
1 Introduction
A Foundation Model (FM) is commonly perceived as a large-scale machine learning model trained on diverse data, capable of being applied to various downstream tasks, which might not necessarily be directly related to its original training objective. The term is coined by The Stanford Institute for Human-Centered Artificial Intelligence’s (HAI) Center in August 2021 as “any model that is trained on broad data (generally using self-supervision at scale) that can be adapted (e.g., fine-tuned) to a wide range of downstream tasks” [1]. Examples of FMs include BERT[2] and the GPT (Generative Pre-Trained Transformer) series[3] in the field of Natural Language Processing (NLP), and CLIP[4] in the field of Computer Vision (CV). Most FMs are constructed based on pre-existing architectures. For example, BERT and the GPT series (language models) are based on the Transformer architecture[5], CLIP (a vision model) is founded on the Vision-Transformer[6], and DriveDreamer[7] (a world model) is built upon the Diffusion Model[8].
1.1 Emergent Abilities
Along with FM, [1] talks about the emergence characteristic or emergent ability of FM as “An ability is emergent if it is not present in smaller models but is present in larger models.” For instance, the adaptability of a language model (LM) to diverse downstream tasks, a novel behavior not directly tied to its initial training, seems to emerge abruptly as the model scales beyond an undisclosed threshold, transforming into a large language model (LLM) [9].
Currently, the emergent abilities of FMs are mainly reflected in Large Language Models (LLMs). In Fig. 1[10] it is illustrated that as the model size, dataset size and the number of computational floats used for training increase, the loss of LLM decreases, providing support for performing large-scale model training. In Fig. 2[9], it can be seen that when the number of parameters of the model reaches a certain level, the performance of the LLMs will get a qualitative leap, showing emergent abilities in different tasks.
Emergent abilities of LLMs are well represented in In Context Learning (ICL)[11]. Context learning ability is a machine learning method to learn in a specific contextual environment. The core idea of this method is to learn from analogies[12], which enables LLMs to get excellent performance in a specific context without parameter tuning.
One particular type of ICL is Chain-of-Thought (CoT)[13]. It allows users to record the derivation process step-by-step for complex problems, such as reasoning problems, and make it available to LLMs. In this way, the LLMs can perform relatively complex reasoning tasks. Emergent abilities are commonly found in FMs, there is currently no compelling explanation for why these abilities would appear the way they do.
Figure 1: Scaling Laws[10]
Figure 2: Emergent abilities of LLMs[9]
Park et al.[14] introduced Generative Agents that simulated real human behaviors, performed daily activities based on pre-input settings, and stored daily memories in natural language. The authors connected Generative Agents to LLM to create a small society with 25 intelligent agents, retrieved memories with LLM, and used its emergent abilities to plan the behaviors of intelligent agents. In the experiment, the intelligent agents emerged with more and more social behaviors in addition to their behaviors, fully demonstrating the LLM’s intelligent emergence.
1.2 Pre-training
The implementation of FMs is based on transfer learning and scaling[1]. The idea of transfer learning[15, 16] is to apply the knowledge learned in one task to another. In deep learning, transfer learning is implemented in two stages, pre-training and fine-tuning. FMs are pre-trained with massive data. After obtaining the pre-trained model a specific dataset is selected for fine-tuning to adapt to different downstream tasks.
Pre-training is the foundation of FMs to obtain emergent abilities. By being pre-trained with massive data, FMs can obtain basic understanding and generative abilities. Pre-training tasks include supervised learning, self-supervised learning, etc[17]. Early pre-training relied on supervised learning, especially in computer vision. To meet the training needs of neural networks, some large-scale supervised datasets, such as ImageNet[18], were built. However, supervised learning also has some drawback, i.e., large-scale data labeling is required. With the gradual increase in the size of the model and the number of parameters, the drawbacks of supervised learning become more obvious. In natural language processing, since the degree of difficulty in labeling text is much greater than that of in labeling images, self-supervised learning is gradually favored by scholars due to its feature of not requiring labeling.
1.2.1 Self-Supervised Learning
Self-supervised learning allows learning feature representations in unlabeled data for subsequent tasks. The distinguishing feature of self-supervised learning is that they do not require manually labeled labels, but instead generate labels automatically from unlabeled data samples.
Self-supervised learning usually involves two main processes[19]: Self-supervised training phase: the model is trained to solve a designed pretext task and automatically generate pseudo labels for the pretext task based on data properties in this phase. The goal is to make the model learn the generic representation of the data.
Downstream tasks application phase: after the completion of self-supervised training, the knowledge learned by the model can be further applied to actual downstream tasks. Downstream tasks use supervised learning methods after self-supervised pre-training of the model. Downstream tasks include semantic segmentation[20], target detection[21], and sentiment analysis[22], etc. Due to the presence of self-supervised training, the generalization ability and convergence speed of the model in downstream tasks will be greatly improved.
Self-supervised learning methods can be generally categorized into three types [23]: generative-based, contrastive-based, and adversarial-based. Generative-based method: first encodes the input data using an encoder and then uses a decoder to regain the original form of the data. The model learns by minimizing the error between the two. And generative-based methods include auto-regressive models, auto-encoding models, etc[24]. Contrastive-based method: constructs positive and negative samples by pretext tasks and learns by comparing the similarity with positive and negative samples. Such methods include SimCLR[25] and others. Adversarial-based method: this method consists of two main components: a generator and a discriminator. The generator is responsible for generating fake samples, while the discriminator is used to distinguish between these fake samples and real samples[23], and a typical example is generative adversarial network[26].
1.2.2 Pretext Tasks of Self-Supervised Learning
The pretext tasks can also be referred to as self-supervised tasks as they rely on the data itself to generate labels. These tasks are designed to enable the model to learn representations that are relevant to a specific task, thereby better handling downstream tasks.
In computer vision, the design of pretext tasks according to data attributes can be summarized into four categories[19]: generation-based, context-based, free semantic label-based, and cross-modal-based. Among them, generation-based approaches mainly involve image or video generation tasks[27, 28]; context-based pretext tasks are mainly designed leveraging contextual features of images or videos, such as contextual similarity, spatial structure, temporal structure, etc.[29, 30, 31]; in the free semantic label-based pretext tasks, the network is trained leveraging automatically generated semantic labels[32]; and cross modal-based pretext tasks need to consider multiple modalities such as vision and voice[33].
In natural language processing, the most common pretext tasks include[34]: center and neighbor word prediction, next and neighbor sentence prediction, autoregressive language modeling, sentence permutation, and masked language modeling, etc. Among them, the Word2Vec[35] model uses center word prediction as a pretext task, while the BERT model uses next sentence prediction and masked language modeling as pretext tasks. These models are trained to learn the expressions of the corpus and applied to downstream tasks.
1.3 Fine-tuning
Fine-tuning refers to further training on a specific task based on an already trained model to adapt it to the specific data and requirements of the task. Typically, a model that has been pre-trained on large-scale data is used as a foundational model, and then it is fine-tuned on a specific task to improve performance. Currently, in the field of LLMs, fine-tuning methods include two main approaches: instruction tuning and alignment tuning[36].
Instruction fine-tuning is aimed to fine-tuning pre-trained models on a collection of datasets described by instructions[37]. Instruction fine-tuning is implemented by collecting or creating instances of instruction formatting and then using these instances to fine-tune the model in a supervised learning manner. Instruction fine-tuning allows LLMs to exhibit strong generalization capabilities on previously unseen tasks.
In most cases, the models obtained after pre-training and fine-tuning can work well, however, some special cases may occur. In the case of LLM, for example, the trained model may appear to fabricate false information or retain biased information from the corpus. To avoid such problems, the concept of human-aligned fine-tuning has been proposed, which aims to make the model’s behavior conform to human expectations[38]. In contrast to instruction fine-tuning, this kind of alignment requires the consideration of completely different standards.
The GPT family is a typical FM, and its training process also includes two stages pre-training and fine-tuning. Taking ChatGPT as an example, the pre-training process of ChatGPT uses self-supervised pre-training [3]. Given an unsupervised corpus, a standard language modeling approach is used to optimize its Maximum Likelihood Estimate(MLE). GPT uses a multi-layer transformer decoder architecture[39], resulting in a pre-trained model.
The fine-tuning phase of ChatGPT can be divided into three steps[38]. In the first step, supervised fine-tuning (SFT) is performed on the obtained pre-trained model; in the second step, comparison data are collected to train the reward model (RM); and in the third step, the SFT model is fine-tuned to maximize the reward leveraging the PPO algorithm[40]. The last two steps together are Reinforcement Learning with Human Feedback (RLHF)[41].
1.4 Applications in Autonomous Driving
The ultimate goal of autonomous driving is to achieve a driving system that can completely replace human driving, and the basic criterion for evaluation is to drive like a human driver. In the past decade, autonomous driving has been rapidly developed in both academia and industry. However, in the face of uncommon or extreme scenario data in autonomous driving scenarios, the existing autonomous driving techniques have limited ability to understand and reason about such scenarios, which still greatly restricts the further development of autonoumous driving. FMs based on pre-training and fine-tuning under the paradigm unleash the great potential of AI in advanced reasoning tasks, and the use of FMs can effectively enhance the understanding and reasoning of the driving scenarios of the autonomous driving system. The research related to the realization of autonomous driving systems by empowering them with FMs has attracted great attention from academia and industry.
In this paper, we provide an overview of the application of FMs in autonomous driving in three segments. Section 2 introduces the application of language and vision FMs in enhancing autonomous driving models. Section 3 introduces the application of language and vision FMs for data augmentation. Section 4 describes the application of World Models, a class of FMs, in exploring the field of autonomous driving. Based on the overview analysis in Sections 2-4, Section 5 presents the challenges and future directions for augmenting autonomous driving with FMs.
2 Autonomous Driving Models
2.1 Human-like Driving Based on Language and Vision Models
With the significant research advances in language models T5[42], GPT-4[43], Llama[4], vision models CLIP[44], ALIGN[45], Segment anything[46], and multimodal models BLIP-2[47], LLaVA[48], and GPT-4V[49] as well as other FMs, its powerful reasoning capability is considered to have brought a new dawn for realizing Artificial General Intelligence (AGI)[50], which has had a significant and far-reaching impact on all aspects of society. In autonomous driving, FMs such as language and vision also show great potential, which is expected to enhance the understanding and reasoning of autonomous driving models on driving scenarios and realize human-like driving for autonomous driving.
In the following, we will provide a review and introduction of the research related to the enhancement of the understanding of the driving scenarios by the autonomous driving system based on FMs of language and vision, as well as the reasoning to give language-guided instructions and driving actions, as illustrated in Fig. 3. Related work on enhancing the understanding of driving scenes is presented in Section 2.1.1, on reasoning to give verbal commands in 2.1.2, and on reasoning to give action commands in 2.1.3.
Figure 3: The pipeline diagram for enhancing autonomous driving leveraging FMs, where FMs refer to language models and vision models. FMs can learn the perceptual information and utilize their powerful ability to understand the driving scenarios and reason to give language-guided instructions and driving actions to enhance autonomous driving.
2.1.1 Understanding of Driving Scenarios
The study by Vasudeva et al.[51] found that the model’s understanding of the scene and object localization can be effectively enhanced by acquiring verbal descriptions and gaze estimation, etc. Li et al.[52] proposed an image captioning model that generates high-level semantic information to enhance the model’s understanding of the traffic scene. Their work verified that linguistic and visual features can effectively enhance the model’s understanding of driving scenes.
Chen et al.[53] transferred image and text features into a 3D point cloud network based on CLIP, which enhanced the model’s understanding of the 3D scenes. Sriram et al.[54] designed and proposed an autonomous navigation framework through a combination of local semantics and natural language commands by fleveraging the natural language coding with the coding of 3D semantic maps obtained from visual images, which was experimentally validated in the CARLA simulator and KITTI dataset[55] for experimental validation. Romero et al.[56] built a video analytics system leveraging a visual-linguistic model, which effectively improves query accuracy based on the powerful comprehension of the visual-linguistic model. Elhafsi et al.[57] improved query accuracy by converting observed visual information into natural language descriptions and passing them to a large language model to identify semantic anomalies leveraging its powerful reasoning ability. Dewangan et al.[58] proposed a method to enhance BEV maps by detecting the features of each object in the BEV through a visual-linguistic multimodal FM (Blip-2[47], Minigpt-4[59], Instructblip[60]) and through linguistic characterizations to obtain a language-enhanced BEV map.
2.1.2 Language-Guided Instructions
Here we present a review of studies that give linguistic commands through FMs, mainly descriptive commands, such as “Red light ahead, you should slow down”, “Intersection ahead, please pay attention to pedestrians”, etc.
Ding et al.[61] used a visual coder to encode video data, which was then fed into a large language model to generate corresponding driving scenario descriptions and suggestions. In particular, this work also proposed a method that enables high-resolution feature maps and the obtained high-resolution information to be fused into multimodal large language models (M-LLMs) to further enhance the model’s recognition, interpretation, and localization capabilities. Fu et al.[62] explored the potential of leveraging LLMs to understand driving environments in a human-like manner, utilizing the LLaMA-Adapter[63] to describe the scene data, and then giving linguistic commands via GPT-3.5. Wen et al.[64] proposed DiLu, a knowledge-driven paradigm based on previous work that can make decisions based on common-sense knowledge and accumulate experience. In particular, the article pointed out that DiLu possesses the ability to direct experience acquisition of real-world data, which has the potential for practical deployment of autonomous driving systems. In the industry, Wayve proposed LINGO-1[65], a grand model of self-driving interaction based on a visual-verbal-action grand model, where the model can interpret itself and answer visually while driving. It introduced human driving experiences, which can explain various causal elements in the driving scenarios through natural language descriptions, acquire feature information in the driving scenario in a human-like understanding way, and learn and give interactive language commands.
2.1.3 Generation of Actions
As described in the previous section 2.1.2, academia and industry have attempted to embed GPT linguistic knowledge into autonomous driving decisions to improve the performance of autonomous driving in the form of linguistic commands to promote the application of large models to autonomous driving. Long before FM made breakthroughs in the LLM field, some works attempted to improve the performance of autonomous driving through similar research ideas. For example, the MP3 framework proposed by Casas et al.[66] used high-level semantic information as a decision training guide, which together with sensory data constitutes the input to build algorithms to realize motion prediction.
The research on the application of LLMs in autonomous driving is on the ascendant, and the GPT series, as the most successful variant of the transformer architecture, may be able to bring breakthroughs to improve the comprehensive performance at multiple levels. LLM is a representative of FMs from the level of linguistic knowledge that empowers the development of autonomous driving, however, linguistic descriptions and reasoning are not directly applied by the autonomous driving system. Considering that the large model is expected to be truly deployed at the vehicle end ,it needs to eventually fall on the planning or control instructions, i.e., FMs should eventually empower the autonomous driving from the action state level. However, how to quantize linguistic decisions into action commands, such as planning and control, available to the autonomous driving system still faces great challenges. Some scholars have already made preliminary explorations, but there is still much room for development. What’s more, some scholars have explored the construction of autonomous driving models through a GPT-like approach, which directly outputs trajectories and even control commands based on LLM. In Table LABEL:tab:1 we provide a brief overview of some representative works.
Sha et al.[67] proposed LanguageMPC, which employs GPT-3.5 as a decision-making component for complex autonomous driving scenarios that require human common sense understanding. By designing cognitive pathways for integrated reasoning in LLM, sha et al. proposed algorithms to transform LLM decisions into actionable driving control commands, which improved the vehicle’s ability to handle complex driving behaviors. Jain et al.[68] achieved navigation localization and further planning of trajectories with the help of visual perception for explicit verbal commands. Omama et al.[69] constructed a multimodal map-based navigation and localization method called ALT-Pilot, which can be used to navigate to arbitrary destinations without the need for high-definition LiDAR maps, demonstrating that off-the-shelf visual language models can be used to construct linguistically enhanced terrain maps.
Some scholars have also attempted to construct autonomous driving models directly through a GPT-like approach, i.e., leveraging LLMs to construct an end-to-end autonomous driving planner, which directly outputs predicted trajectories, path planning, and even control commands, intending to effectively improve the ability of autonomous driving models to generalize to unknown driving scenarios. Keysan et al.[70] took information about the current surroundings on the road as input and then output a sequence that includes future vehicle decision instructions and spatial paths that can be followed by the self-driving vehicle. Seff et al.[71] proposed MotionLM, which uses motion prediction as a language modeling task to learn multimodal distributions by representing continuous trajectories as discrete sequences of motion tokens leveraging a single standard language modeling objective to predict the future behaviors of road network participants. Mao et al.[72] proposed the GPT-Driver model to reformulate motion planning as a language modeling problem by representing the inputs and outputs of the planner as linguistic tokens and leveraging the LLM to generate driving trajectories through linguistic descriptions of coordinate positions. Furthermore, the model setup allows for the description of internal decision-making processes in natural language, which also improves the interpretability of the decisions. Pallagani et al.[73] constructed Plansformer, which is both a large language model and a planner, showing the great potential of a large language model fine-tuned as a planner from a variety of planning tasks. Wang et al.[74] constructed the BEVGPT model, which takes as input information about the current surroundings on the road and then outputs a sequence that includes future vehicle decision instructions and spatial paths that can be followed by self-driving vehicles. DriveGPT4 proposed by Xu et al.[75] is further used for autonomous driving planning control by inputting video data, textual data, and historical control signals and outputs as textual responses to the problem as well as predictions of control signals to be used for vehicle operation. Cui et al.[76] utilized GPT-4 with inputs of natural language descriptions and environment perception data to make LLM directly output driving decisions and operation commands, and further experimented with highway overtaking and lane changing scenarios in Ref.[77] to compare driving decisions provided by LLM with different cues, and the study showed that chained-thinking cueing helps LLM to make better driving decisions. Considering the scale challenges of Visual Language Model (VLM), Chen et al.[78], based on the idea that digital vector modalities are more compact than image data, fused vectorized 2D scene representations with pre-trained LLMs to improve the LLM’s ability to interpret and reason about the integrated driving situation, giving scene interpretation and vehicle control commands. Mao et al.[mao2023language] proposed Agent Driver, which utilized LLM to introduce a general-purpose library of tools accessible via function calls, cognitive memory for common sense and empirical knowledge for decision making, and a reasoning machine capable of chain-of-thinking (CoT) reasoning, task planning, motion planning, and self-reflection to achieve a more nuanced, human-like approach to autopilot.
This type of research idea is much closer to human driving than pure knowledge embedding to make an autonomous driving model and perhaps has the potential to become one of the main development directions in the future with the development of large models. However, it should be noted that these new frameworks are also questionable in terms of reliability due to the unresolved pitfalls of large models themselves, such as “illusions” (LLMs may generate content that conflicts with source or factual information). Specific details about the problems of the large models themselves and the challenges inherited in autonomous driving will be discussed in detail in Section 5.
Table 1: Works on the use of LLMs for generating autonomous driving planning and control
Authors Input Output Learning Description
Sha et al.[67] 2023 Prompt of scenario Control Action Sequences SL This work enables navigation and localization as well as further trajectory planning through visual perception and verbal commands.
Omama et al.[69] 2023 OSM Maps with Descriptions, LiDAR, Camera Location, Trajectory SL This work utilizes a Bayesian state estimation model fleveraging visual-linguistic features to generate global paths, plan trajectories, and control vehicles to complete navigation.
Keysan et al.[70] 2023 Scene Raster, Text Prompt Trajectory SL This work encodes the rasterized scene and the text prompt with pre-trained models dedicated for each modality, finally picks the target trajectory from a pre-generated trajectory set.
Seff et al.[71] 2023 Multimodal Scene Motion Prediction SL This work uses a single standard linguistic modeling objective to learn multimodal distributions for predicting the future behavior of traffic participants.
Mao et al.[72] 2023 Perception, Ego-States, Trajectory, Goal Thoughts, Driving Decisions, Trajectory SL This work represents the inputs and outputs as linguistic tokens and utilizes the LLM to generate driving trajectories and provide explanations for decision making.
Wang et al.[74] 2023 Scene Information Driving Decisions, Trajectory SL In pre-training stage, this work trains a causal transformer for driving scenario prediction and decision-making. In fine-tuning stage, it adapts to motion planning and accurate BEV generation.
Xu et al.[75] 2023 Vedio, Text Prompt, Control Signal Decision, Control Signal SL Video sequences undergo tokenization leveraging a tokenizer, while text and control signals share a common tokenizer. Following tokenization, the model can generate responses to human inquiries and predict control signals.
$\bullet$ This “SL” refers to “Supervised Learning”.
2.2 End-to-end Autonomous Driving with Pretraining Backbone
As mentioned in 2.1, LLMs-based autonomous driving-related research mainly focuses on improving the interpretability of autonomous driving decisions, enhancing scenario understanding, and guiding the planning or control of autonomous driving systems. The research idea of “pre-training + fine-tuning” in autonomous driving research does not only appear after the introduction of large models but has been researched for a long time. To use a more familiar term, it is end-to-end autonomous driving. In the past few years, scholars have been optimizing Pretraining Backbone in various ways, including the transformer architecture and self-supervised learning methods——Note that we define Pretraining Backbone here as a model that transforms each modal input into a usable feature representation for downstream tasks (such as target detection, trajectory prediction, decision planning, etc.). There have also been many research attempts to develop end-to-end frameworks based on the transformer architecture, with excellent results. Hence, to summarize the application of the underlying models in autonomous driving more comprehensively, we believe that it is necessary to introduce the research related to end-to-end autonomous driving based on Pretraining Backbone. In this section, we review the latest research on Pretraining Backbone with end-to-end autonomous driving solutions, the pipeline of such methods is briefly illustrated in Fig. 4.
Figure 4: The pipeline diagram for the end-to-end autonomous driving system with pretraining backbone. Multimodal sensing information is input to pretraining backbone to extract features, after which it enters into the framework of autonomous driving algorithms built by various methods to realize tasks, such as planning/control, to achieve end-to-end autonomous driving.
2.2.1 Pretraining Backbone
In the end-to-end modeling, feature extraction of low-level information from raw data determines the potential of subsequent model performance to a certain extent, and an excellent Pretraining Backbone can endow the model with more powerful feature learning capability.
Pretraining convolutional networks such as ResNet[79], and VGG[80] are the most widely used backbone for visual feature extraction in end-to-end models. These pretraining networks are often trained leveraging target detection or segmentation as the task to extract generalized feature information, and their competitive performances have been verified in many works. ViT[6] first applied the transformer architecture to image processing and achieved excellent classification results. Transformer has the advantage of optimized algorithms for handling large-scale data with its simpler architecture and faster inference speed. The self-Attention mechanism is very suitable for processing time series data, there is the possibility of modeling and predicting the temporal motion trajectories of objects in the environment, and is conducive to the fusion of heterogeneous data from multiple sources, such as LiDAR point clouds, images, maps, etc.
Another class of Pretraining Backbone, represented by LSS[81], BEVDet[82], BEVformer[83], BEVerse[84], etc., expanded the usability by extracting the images captured by the surround-view camera and converting them to bird’s-eye view (BEV) features through model learning, indexing the local image features from the two-dimensional viewpoint to the three-dimensional space. In recent years, BEV has attracted extensive attention due to its ability to describe the driving scene more accurately, and the research of leveraging BEV features such as Pretraining Backbone output is not limited to camera, and the extraction and fusion of multimodal sensing BEV features represented by BEVfusion[85] has further provided a wider vision for autonomous driving systems. However, it should be pointed out that although the transformer architecture brings great performance enhancement, these Backbone still constructs pre-trained models with supervised learning, relying on massive labeled data, and the quality of the dataset also greatly affects the final result of the model.
In both camera and point cloud processing domains, some works implement pre-trained Backbone by unsupervised or self-supervised learning. Wu et al.[86] proposed the PPGeo model, which uses a large number of unlabeled driving videos to accomplish the pre-training of the visual coder in two stages, and can be adapted to different downstream end-to-end autonomous driving tasks. Sautier et al.[87] proposed BEVContrast for self-supervision of 3D Backbone on automotive LiDAR point clouds, which defines contrasts at the level of 2D Cells in the BEV plane, retaining the simplicity as in PointContrast[88] while maintaining good performance in downstream driving tasks.
2.2.2 End-to-end Autonomous Driving Models
Early work on end-to-end autonomous driving modeling was mainly based on various types of deep neural networks, which were constructed through imitation learning[89, 90, 91, 92] or reinforcement learning[93, 94, 95] methods. The work of Chen et al.[96] analyzed the key challenges facing end-to-end autonomous driving from a methodological perspective, pointing out the future trend of empowering end-to-end autonomous driving with fundamental models such as transformer. Some scholars have tried to build an end-to-end autonomous driving system with transformer and got competitive results. For instance, there have been Transfuser[97, 98], NEAT[99], Scene Transformer[100], PlanT[101], Gatformer[102], FusionAD[103], UniAD[104] and a host of end-to-end frameworks developed based on transformer architecture.
Chitta et al.[97, 98] proposed Transfuser, which takes RGB images and BEV views from LiDAR as inputs, uses multiple transformers to fuse the feature maps, and predicts the trajectory points for the next four steps through a single-layer GRU network, followed by longitudinal and transverse PIDs to control the vehicle operation. NEAT[99] further mapped the locations of BEV scene coordinates to trajectory points and semantics and compresses high-dimensional image features leveraging an intermediate attention map, which allows the model to selectively focus on driving-relevant regions and ignore driving task-irrelevant information. PlanT proposed by Renz et al.[101] used simple object-level representations (vehicles, roads) as inputs to the transformer encoder, and used speed prediction of surrounding vehicles as a secondary task to predict future waypoint trajectories. UniAD proposed by Hu et al.[104] enhanced the design of the decoder and achieving the integration of the full stack of autonomous driving tasks into a unified framework to improve autonomous driving performance, although still relying on different sub-networks for each task. This work also won the CVPR 2023 Best Paper Award, which shows the academic recognition of the end-to-end autonomous driving paradigm.
Moreover, Drive Anywhere proposed by Wang et al.[105] not only realize end-to-end multimodal autonomous driving but also combine with LLM to be able to provide driving decisions based on representations that can be queried through images and texts. Dong et al.[106] generated image-based action commands and explanations by building a feature extraction model based on transformer. Jin et al.[107] proposed the ADAPT model to directly output vehicle control signals with inference language descriptions through an end-to-end model. This is the first driving action captioning architecture based on an action-aware transformer. It accomplished the driving control task while adding natural language narratives to guide the decision-making and action process of the autonomous driving control module. It also helped the user to get the vehicle’s state and the surrounding environment at all times, and to batter understand the basis of the actions taken by the autonomous driving system, which improved the interpretability of the decision-making. This provides a glimpse of the potential of the transformer architecture to enhance the interpretability of end-to-end driving decisions.
3 Data Augmentation
With the further development of deep learning, the performance of FMs with pre-training and fine-tuning as the underlying architecture is getting better and better, and FMs represented by GPT are leading the learning paradigm to shift from rule-driven to data-driven. The significance of data as a key link in model learning is self-evident. In the training and testing process of autonomous driving models, a large amount of scenario data is used to enable the models to have a good understanding and decision-making ability for various road and traffic scenarios. The long-tail problem faced by autonomous driving is also the very fact that there are endless unknown edge scenarios that make the generalization ability of the model seem perpetually insufficient, leading to poor performance. In Table LABEL:tab:2 we provide a brief overview of some representative works.
Data augmentation is crucial for improving the generalization ability of autonomous driving models. The realization of data augmentation needs to consider two aspects: on the one hand, how to obtain large-scale data so that the data fed to the autonomous driving model is diverse and extensive; on the other hand, how to obtain as much high-quality data as possible so that the data used to train and test the autonomous driving model is accurate and reliable. Related research have also roughly chosen two directions to enhance the autonomous driving data, one is to enrich the data content of existing datasets and enhance the data features of driving scenarios, and the other is to generate driving scenarios with multiple levels through simulation. In the following, a review of related research on enhancing data based on language and vision FMs will be presented, in section 3.1 we describe related work on extending datasets, and in 3.2 we describe related work on generating driving scenarios.
3.1 Expansion of Autonomous Driving Datasets
Existing autonomous driving datasets are mostly obtained by recording sensor data and then labeling the data. The features of the data obtained in this are usually low-level and exist more at the level of numerical representation, which is insufficient for the visuo-spatial features of the autonomous driving scenarios. Natural language descriptions are seen as an effective way to enhance scene representation[4], Flickr30k[108], RefCOCO[109], RefCOCOg[110], and CLEVR-Ref[111] use concise natural language descriptions to identify the corresponding visual regions in an image. Talk2Car[112] fused image, radar, and LiDAR data to construct the first object-referenced dataset containing natural language commands for self-driving cars. However, the Talk2Car dataset allowed only one object to be referenced at a time. CityFlow-NL[113] constructed a dataset for multi-target tracking through natural language descriptions, and Refer-KITTI[114] achieved prediction of arbitrary target tracking by leveraging natural language queries in the corresponding task.
FMs represented by language models provide new ideas for enriching and expanding autonomous driving datasets under their advanced semantic understanding, reasoning, and interpretation capabilities. Qian et al.[115] created NuScenes-QA, a visual question and answer dataset for autonomous driving in 3D multi-view driving scenarios, by encoding question descriptions through a language model and obtaining answers through feature fusion with sensor data. Significant progress was made in the use of natural language prompts. Wu et al.[116] extended NuScenes-QA by constructing the dataset NuPrompt by capturing and combining natural language elements and then invoking LLM to generate the descriptions. The dataset provided a finer match between the 3D instances and each of the prompts, which helped to more accurately characterize objects in the autopilot images, and improved the accuracy of the model. In addition to directly extending the augmented autonomous dataset, some scholars have also integrated the chain-of-thinking capability of LLM and the cross-modal capability of the vision model to build an automatic annotation system, OpenAnnotate3D[117], which can be used for multimodal 3D data. Expanding the dataset by utilizing the advanced understanding, reasoning, and interpretation capabilities of the underlying models can help to better assess the interpretability and control of the autonomous driving system, thus improving the safety and reliability of the autonomous driving system.
3.2 Generation of Driving Scenarios
Driving scenarios are of great significance for autonomous driving, and to obtain diverse driving scenario data, real-time acquisition of vehicle sensors alone requires huge costs, and it is difficult to obtain sufficient scenario data for some edge scenarios. Generating realistic driving scenarios through simulation has attracted the attention of many researchers, and traffic simulation research can be mainly divided into rule-based and data-driven. Rule-based approaches[118, 119, 120, 121] require the use of predefined rules, which are generally insufficient for the characterization of complex driving scenarios, and the simulated driving scenarios are simpler and less generalizable. Data-driven approaches[122, 123, 124, 125] refer to the use of driving data to train models, from which the models can continuously learn and adapt. However, data-driven approaches often require a large amount of labeled data for training, which hinders the further development of traffic simulation. The current emergence of some high-quality datasets KITTI[126], APolloScape[127], and nuScenes[128] has somewhat remedied this limitation. However, how to construct an effective learning paradigm to get rid of data labeling becomes more important. Recently, great success has been achieved with generative large models obtained by training based on natural language data, and such models can utilize self-supervised learning to learn a wide range of data.
Enhancing data-driven approaches through generative FMs, such as language and vision models, to accelerate traffic simulation has attracted significant research attention. Marathe et al.[129] efficiently generated a dataset comprising 16 weather extremes by constructing a cueing paradigm leveraging a visual-linguistic model. However, the model had some extension constraints due to the phenomenon of pre-selected fixation in data selection. In response to the fact that some long-tailed scenes can never be collected in multi-view shots, Yang et al.[130] fused verbal cues, BEV sketch, and multi-view noise to design a two-stage generative network BEVControl for synthesizing realistic street scene images. However, BEVControl is insufficient for modeling foreground and background detail information. To address the difficulty of obtaining large-scale BEV representations, Li et al.[131] proposed a spatiotemporally consistent diffusion framework, DrivingDiffusion, to autoregressively generate realistic multiview videos controlled by 3D layouts. The quality of the generated data can be effectively improved by introducing local cue inputs into the vision model. Based on the powerful language understanding, reasoning, and interpretation capabilities of LLM, it has also become a research hotspot to combine LLM into driving scenario generation. Chen et al.[132] realized the combination of numeric vector modality and natural language by pairing control commands collected by reinforcement learning intelligence and question answers generated by LLM to directly construct new data. Zhong et al.[133] proposed a scenario-level diffusion-based language-guided traffic simulation model, CTG++, which can generate instruction-compliant, realistic, and controllable traffic scenarios. Wang et al.[105] utilized natural language descriptions as conceptual representations that were integrated with LLM to enrich the complexity of the generated scenarios leveraging their powerful common-sense reasoning capabilities. The behavior of human drivers is also an important part of driving scenarios, Jin et al.[134] proposed SurrealDriver, a generative driving agent simulation framework for urban environments based on LLM. By analyzing and learning from real driving data, SurrealDriver can capture the driver’s behavioral patterns and decision-making processes and generate behavioral sequences that are similar to those in real driving. similar behavioral sequences as in real driving.
Through efficient and accurate simulation of driving scenarios, including a variety of complex and dangerous situations, a large amount of training data is provided for model learning, which can effectively enhance the generalization capability of autonomous driving systems. Meanwhile, the generated driving scenarios can also be used to evaluate different autonomous driving systems and algorithms to test and verify their system performance.
Table 2: Works on Data Augmentation
Authors Input Output Learning
Expand Dataset Qian et al.[115] 2023 Images, Text, Point Clouds Q-A Pair SL
Wu et al.[116] 2023 Images, Text Object-Text Pairs SL
Zhou et al.[117] 2023 Images, Text Labeled Data SL
Generate Scenarios Marathe et al.[129] 2023 Objects, Scenes, Weather Condition Multi-Weather Images SL
Yang et al.[130] 2023 Text, BEV Sketch, Multi-View Noise Street-View Images SL
Li et al.[131] 2023 Layouts, Frames, Optical Flow Prior Multi-View Videos SL
Chen et al.[132] 2023 Objects, Text Trajectory SL
Zhong et al.[133] 2023 Text, Scenes with Noise Traffic Scenes SL
Wang et al.[105] 2023 Images, Knowledge Latent Space Simulation SL
Jin et al.[134] 2023 Text, Simulation of Urban Driving Driving Maneuvers SL
$\bullet$ This “SL” refers to “Supervised Learning” and “Q-A” refers to “Question-Answer”.
4 World Model
World Models refer to the mental models of the world. It can be interpreted as a type of artificial intelligence model that encompasses a holistic understanding or representation of the environment in which it operates. This model is capable of simulating the environment to make predictions or decisions. The term “World Models” has been mentioned in the context of reinforcement learning in recent literature[137, levine2021understanding]. This concept has also gained attention in autonomous driving applications due to its capacity to comprehend and articulate the dynamics of the driving environment, as will be detailed below. In his position paper, LeCun[135] pointed out that the learning capability of humans and animals may be rooted in their capacity to learn World Models, allowing them to internalize and understand how the world works. He pointed out that humans and animals have exhibited the ability to acquire a vast amount of background knowledge about the functioning of the world observing a small number of events, whether related or unrelated to the task at hand.
The idea of World Model can be traced back as far as Dyna, proposed by Sutton et al.[136] in 1991, where observing the state of the world and selecting an action based on its response learns interactively with the world. Dyna is essentially a form of reinforcement learning under supervised conditions. After that, researchers have also made many attempts. Ha et al.[137] attempted to learn leveraging an unsupervised approach, VAE to encode input features and RNN to learn the evolution of the state. Hafner et al.[138] proposed the Recurrent state space model (RSSM), which combined reinforcement learning to realize multi-step prediction that integrates stochasticity and determinism. Based on the RSSM architecture, Hafner et al. successively proposed DreamerV1[139], DreamerV2[140], and DreamerV3[141], which learned in implicit variables to realize image prediction generation. Gao et al.[142] considered that there was redundant information implicit, and extended the framework of the Dreamer series by proposing the Semantic Masked recurrent World Model (SEM2) to learn relevant driving states. Hu et al.[143] removed prediction rewards and proposed an imitation learning-based method MILE to predict future states.
It can be seen that World Model is highly related to reinforcement learning, imitation learning, and deep generative models. However, utilizing World Models in reinforcement learning and imitation learning generally requires labeled data, and both SEM2 and MILE approaches mentioned are conducted within a supervised paradigm. There have also been attempts to combine reinforcement learning and unsupervised learning based on the limitations of labeled data[144, 145]. Due to the close relationship with self-supervised learning, deep generative models have become more and more popular, and many attempts have been made by researchers in this field. In the following, we will mainly review the exploratory applications of generative World Models in autonomous driving, the pipeline is illustrated in Fig. 5, 4.1 introduces the principles of various types of deep generative models and their applications in generative driving scenarios, 4.2 introduces the applications of generative World Models in autonomous driving, and 4.3 will introduce a class of non-generative methods.
Figure 5: The pipeline diagram for enhancing autonomous driving with World Models. The World Models first learn the intrinsic evolutionary patterns by observing the traffic environment and then enhance the autonomous driving by hooking up different decoders adapted to different driving tasks.
4.1 Deep Generative Models
Deep generative models generally include variational autocoders (VAEs) [146, 147], generative adversarial networks (GANs)[26, 148], flow models[149, 150], and autoregressive models (ARs)[8, 151, 152].
VAEs combine the ideas of self-encoders and probabilistic graphical models to learn the underlying structure of the data and generate new samples. Rempe et al.[153] used VAE to learn prior distributions of traffic scenarios and simulate the generation of accident-prone scenarios. GANs consist of two neural networks, the Generator and the Discriminator, which compete and enhance each other utilizing adversarial training, to ultimately achieve the goal of generating realistic samples. Kim et al.[154] used a GAN model to observe sequences of unlabeled video frames and their associated action pairs to simulate a dynamic traffic environment. The flow models generate similar data samples by transforming simple prior distributions into complex posterior distributions through a series of invertible transformations. Kumar et al.[155] used the flow model to achieve multi-frame video prediction. ARs are a class of sequence analysis methods, based on the autocorrelation between the sequence data, describing the relationship between the current observations and the past observations, and the estimation of the model parameters is usually done leveraging the least squares method and maximum likelihood estimation. For example, GPT uses maximum likelihood estimation for model parameter training. Feng et al.[156] achieved the generation of future trajectories of vehicles based on autoregressive iterations. The diffusion Model is a typical autoregressive model that learns the process of gradual denoising from purely noisy data. With its strong generative performance, the Diffusion Model is the new SOTA among current deep generative models. Works such as[157, 158, 159] demonstrated that the Diffusion Model has a strong ability to understand complex scenes, and the video diffusion model can generate higher quality videos. Work such as[160, 161] utilized the Diffusion Model to generate complex and diverse driving scenarios.
4.2 Generative Methods
Based on the powerful capabilities of deep generative models, leveraging deep generative models as World Models to learn driving scenarios to enhance autonomous driving has become popular. The following section will review the applications of leveraging deep generative models as World Models in autonomous driving. In Table LABEL:tab:3 we provide a brief overview of some representative works.
Karlsson et al.[162] used a hierarchical VAE to construct World Model, used latent variable prediction and adversarial modeling to generate pseudo-complete states, matched partial observations with pseudo-complete observations to predict future states, and evaluated it on the KITTI-360[163] dataset. In particular, it utilized pre-trained vision-based semantic segmentation models to infer semantic information from image observations without manual annotation. Similarly, to address the challenges of predicting future changes in driving scenarios, Wayve proposed a generative World Model, GAIA-1[7]. GAIA-1 used transformer as World Model, learned and predicted the next states of the input video, text, and action signals and then generated a realistic driving scenarios. For the learning of video streams, GAIA-1 adopted self-supervised learning, which can learn scaled data and obtain a comprehensive understanding of driving scenarios. Wang et al[164] designed a two-phase training strategy that incorporates the features of the diffusion model to learn the driving scenarios and obtain an understanding of the structured traffic information. The driving World Model is then modeled through video prediction task, specifically, the prediction is formalized into a generative probabilistic model to generate future driving scenarios. Moreover, by integrating historical driving behavior, this method can also generate future driving actions. zhang et al.[165] built on Maskgit[166] and recast it into a discrete diffusion model for point clouds prediction. This method utilized VQ-VAE[167] to tokenize the observation data for label-free learning. Bogdoll et al[bogdoll2023muvo] constructed a multimodal autonomous generative World Model, MUVO, leveraging raw images and LiDAR data to learn a geometric representation of the world. And conditioned on actions, this mode achieved 3D occupancy prediction and can be directly used for downstream tasks(e.g.,planning). Similarly, Min et al.[min2023uniworld] used unlabeled image-LiDAR for pre-training to construct World Model that can generate 4D geometric occupancy. This work intends to obtain a FM through pre-training and subsequently fine-tuning on downstream tasks. In the industry, at the 2023 CVPR Autonomous Driving Workshop, Tesla researcher Ashok presented exploratory work in utilizing generative large model[168] to generate future driving scenarios. And in the demonstration it was seen that the videos generated by Tesla’s generative large model were very close to those captured from real vehicles. It also can generate annotation-like semantic information, indicating that the model also has some semantic-level understanding and reasoning capabilities. Tesla named their work “Learning a General World Model” and it can be seen that their understanding is to build a generalized World Model. By learning from a large amount of video data captured from real vehicles, Tesla intends to build a large-scale FM for autonomous driving, which can understand the dynamic evolution of the world by learning from image data at scale.
Vision is one of the most direct and effective ways for humans to obtain information about the world because the feature information contained in image data is extremely rich. Numerous previous works[139, 140, 141, 145, 169] have accomplished the task of image generation through World Model, demonstrating that World Model has a good understanding and reasoning ability for image data. However, these are mainly focused on image generation and are still lacking in video prediction tasks that can better represent the dynamic evolution of the world. Video prediction tasks require a deeper understanding of world evolution and also stronger guidance for downstream tasks. In the research works[7, 162], they all effectively predicted generated future traffic scenarios, where self-supervised learning may be key. Previous work has explored this as well. Villegas et al.[170] trained a model leveraging raw images and proposed a hierarchical long-term video prediction method combining low-level pixel space and high-level feature space (e.g., landmarks), achieving longer video prediction compared to the work[141]. Endo et al.[171] constructed a model under the self-supervised learning paradigm for predicting future traffic scenes from single-frame images to predict the future. Based on a denoising diffusion model with probabilistic conditional scores, Voleti et al.[172] trained the model by randomly masking the past frames or future frames unlabeled, which allowed block-by-block autoregressions to generate videos of arbitrary length. Finn et al.[173] proposed to physically interact with the world under unsupervised conditions and realized video prediction by predicting the distribution of pixel motion in the previous frame. Micheli et al.[174] verified the effectiveness of leveraging an autoregressive transformer as a World Model, and achieved the prediction of game images by training the parameters through self-supervised learning. Wu et al.[175] constructed an object-centered World Model to learn complex spatio-temporal interactions between objects and generated high visual quality future prediction.
From the above research, we can see that researchers hope to learn the intrinsic evolutionary laws of the world from image data, and then predict the future state. Combined with self-supervised learning, the World Model is used to learn from image data, fully unleashing the inference ability of the model, and constructing a generalised Foundation Model for the visual domain may be promising.
Table 3: Works on the use of World Models for prediction
Authors Input Output Learning Description
AD Karlsson et al.[162] 2023 Images, Point Clouds Semantic Point Clouds SSL This work utilizes a hierarchical VAE to construct World Model and generates pseudo-complete states and matches them with partial observations to predict future states.
Hu et al.[7] 2023 Videos, Text, Actions Videos SSL This work utilizes an autoregressive transformer to construct World Model and leverages DINO, a self-supervised image model, to tokenize images.
AD Wang et al.[164] 2023 Images, HDMap, 3D Box, Text, Actions Videos, Actions SSL This work obtains comprehension of the structured traffic information. Then the prediction is formalized into a generative probabilistic model.
Zhang et al.[165] 2023 Point Clouds, Actions Point Clouds UL/SSL This work utilizes a discrete diffusion model for point cloud prediction, which is a spatio-temporal transformer. And this work leverages VQ-VAE to tokenize sensor observations.
Min et al.[min2023uniworld] 2023 Image-LiDAR pairs 4D GO UL/SSL This work proposes a spatial-temporal World Model for unified autonomous driving pretraining.
Bogdoll et al.[bogdoll2023muvo] 2023 Actions, Point Clouds, Images Point Clouds, Images, 3D OG UL/SSL This work leverages raw data to learn a sensor-agnostic 3D occupancy representation and predicts future states conditional on actions.
VP Finn et al.[173] 2016 Videos Videos UL/SSL This work develops an action-conditioned video prediction model for interaction that incorporates appearance information in previous frames with motion predicted by the model.
Micheli et al.[174] 2022 Images, Actions Images SSL The World Model is composed of a discrete autoencoder and an autoregressive transformer. And it is trained in a self-supervised manner.
Wu et al.[175] 2022 Videos Videos UL/SSL This work leverages a pre-trained object-centric model to extract object slots from each frame. These slots are then forwarded to a transformer and used to predict future slots.
$\bullet$ This “AD” refers to “Autonomous Driving”and this “VP” refers to “Visual Prediction”. $\bullet$ This “GO” refers to “Geometric Occupancy” and this “OG” refers to “Occupancy Grids”. $\bullet$ This “SSL” refers to “Self-Supervised Learning” and this “UL” refers to “Unsupervised Learning”.
4.3 Non-Generative Methods
In contrast to generative World Models, LeCun[135] elaborated his different conception of World Model by proposing a Joint Extraction and Prediction Architecture (JEPA) based on Energy-based Model. This is a non-generative and self-supervised architecture, as it does not predict the output $y$ directly from the input $x$, but encodes $x$ as $sx$ to predict $sy$ in an abstract space. This has the advantage that it does not have to predict all the information about $y$, and can eliminate irrelevant details.
Since its proposal, the JEPA architecture has been applied by several scholars in different fields with excellent performance. In computer vision, Assran et al.[176] proposed I-JEPA for learning highly semantic image representations without relying on manual data enhancement. Combining I-JEPA with ViT resulted in strong downstream performance in a variety of tasks ranging from linear classification to object counting and depth prediction. Bardes et al. proposed MC-JEPA in Rrf.[177], which uses the JEPA architecture and a self-supervised learning approach to achieve co-learning of optical flow and content features from video to learning dynamic content features. MC-JEPA performs well in various tasks such as optical flow estimation, and segmentation of images and videos. In the graph domain, Skenderi et al. proposed Graph-JEPA[178], which is a JEPA model for graph domains. It divides the input graph into subgraphs and then predicts the representation of the target subgraph in the context subgraph. Graph-JEPA has obtained excellent performance in both graph classification and regression problems.
JEPA is a self-supervised learning architecture based on Energy-based Model that observes and learns the laws of how the world works and is highly generalizable. There is also great potential for JEPA in autonomous driving, which is expected to generate high-quality driving scenarios and driving strategies by learning how driving works.
5 Conclusion and Future Directions
This paper provides a comprehensive overview of FMs in autonomous driving. Recent works on the application of FMs, notably language and vision models, towards advancing autonomous driving are summarized in detail in Section 2, and their involvement in shaping the research of end-to-end autonomous driving is extensively surveyed. Recent works on data augmentation of FMs are detailed in Section 3. In Section 4 we present an exploratory application of the World Model in autonomous driving. Overall, FMs can effectively contribute to autonomous driving in terms of both data enhancement and model optimization.
However, many of the aforementioned studies also raised concerns about the maturity of FMs. Their suitability for the safety-critical downstream task of autonomous driving requires further development and is awaiting confirmation. Whether it is the hallucinations[zhang2023siren, liu2023exposing], security concerns such as inference latency[pope2023efficiently, weng2023inference], limited context length[liu2023lost, longchat2023], or practical issues such as storage limitations, deployment costs, and ethical and legal constraints that have yet to be established, the solution of “autonomous driving through foundation models” to achieve autonomous driving on public roads across various weather and traffic conditions still requires a comprehensive and holistic approach. Among them, the hallucination error problem is mainly manifested as misrecognition in autonomous driving, such as wrong target detection, which may cause serious safety accidents. The phantom problem mainly arises because of the limited samples in the dataset or because the model is affected by unbalanced or noisy data, and the stability and generalization ability of the model needs to be strengthened by means of expanding data and adversarial training. Challenges in security are also severe, while inference latency may result in vehicles not being able to make timely decisions and execute them, posing serious traffic safety risks. FMs have seen an amazing emergent ability with model scaling, and demonstrated success in natural language processing. However, in the context of autonomous driving, this line of research faces additional open challenges due to limited available data and extended context length issues. These challenges contribute to an inadequate understanding of macroscopic driving scenarios, thereby complicating long-term planning in this field. Driving video is a high-dimensional continuous modal with an extremely large amount of data (several orders of magnitude larger compared to textual data). Hence, training large visual model requires a more macroscopic scene distribution to embed enough video frames to reason about complex dynamic scenes, which requires a more robust network structure and training strategies to learn this information. In a recent study, Bai et al.[bai2023sequential] proposed a two-stage approach, in which images are converted into discrete tokens to obtain “visual sentences”, and then autoregressive predictions are made, similar to the standard approach for language models[11]. Another promising solution may lie in World Models. As described in Section 4, World Models can learn the intrinsic evolutionary laws of the world by observing a small number of events that are either relevant or irrelevant to the task. However, World Models also have certain limitations in exploratory applications, where uncertainty in the predictive outcomes of the models, as well as learning what kind of data captures the intrinsic laws of how the world works, still warrant further exploration. Multi-modal data fusion is expected to provide more feature information and may be a promising research direction.
From the findings of related research, another direction worth further exploration is how to design an effective architecture for self-supervised learning. Self-supervised learning can effectively break through the limitation of data labeling, allowing models to learn data at scale and fully unleash the inference capability of the model. If a Foundation Model for autonomous driving can be trained using scaled driving scenario data under the paradigm of self-supervised learning, it is anticipated that its generalization and emergent capability will be greatly improved. This advancement may lead to the realization of a more generalized Foundation Model. In conclusion, although there are many challenges to be solved in applying Foundation Models to autonomous driving, its potential has already begun to take shape. In future, we will continue to monitor the progress of Foundation Models applied to autonomous driving.
Acknowledgments
Author Contributions
Hong Chen, Hongqing Chu, B.G. and J.W. conceived and designed the study. J.W., J.G. and J.Y. wrote the manuscript draft. X.C., Y.C., Q.Y. and H.E.T gave some helpful suggestions. Hong Chen, Hongqing Chu, B.G. and J.C. revised the manuscript. The authors read and approved the final manuscript.
Funding
This work was supported in part by the National Nature Science Foundation of China (No. 62373289, 62273256), and the Fundamental Research Funds for the Central Universities.
Conflicts of Interest
The authors declare that there is no conflict of interest regarding the publication of this article.
[…]